OpenStudy (anonymous):

-[-5m-3(m-n)+7]-(2m+n)

OpenStudy (anonymous):

\(-[-5m-3(m-n)+7]-(2m+n)=-[-5m-3m+3n+7]-2m-n\) \(=8m-3n-7-2m-n=6m-4n-7\)

jhonyy9 (jhonyy9):

dear angela your answer not is correct

OpenStudy (anonymous):

tnx

OpenStudy (angela210793):

:/